solve (x+5)^2=36 thanks
Answer by checkley71(8403)

(X+5)^2=36
X62+10X+25=36
X^2+10X+25-36=0
X^2+10X-9=0 USING THE QUADRATIC EQUATION WE GWT
X=(-B+-SQRT[B^2-4AC])/2A
X=(-10+-SQRT[100-4*1*-9])/2*1
X=(-10+-SQRT[100+36])/2
X=(-10+-SQRT[136])/2
X=(-10+-11.66)/2
X=(-10+11.66)/2
X=1.66/2
X=.83 SOLUTION
X=(-10-11.66)/2
X=-21.66/2
X=10.83 SOLUTION